# Rheo-SINDy: Finding a constitutive model from rheological data for complex fluids using sparse identification for nonlinear dynamics

### v2

@@ Abstract
Rheology plays a pivotal role in understanding the flow behavior of fluids by discovering governing equations that relate deformation and stress, known as constitutive equations. Despite the importance of these equations, current methods for deriving them lack a systematic methodology, often relying on expert knowledge and incurring substantial costs. To overcome this problem, we propose a novel method named Rheo-SINDy, which employs the sparse identification of nonlinear dynamics (SINDy) algorithm for discovering constitutive models from rheological data. Rheo-SINDy was applied to five distinct scenarios, four with well-established constitutive equations and one without predefined equations. Our results demonstrate that Rheo-SINDy successfully identified accurate models for the known constitutive equations and derived physically plausible approximate models for the scenario without established equations. Notably, the identified approximate models can accurately reproduce nonlinear shear rheological properties, including shear thinning. These findings validate the robustness of Rheo-SINDy in handling real-world data complexities and underscore its efficacy as a tool for advancing the development of data-driven approaches in rheology.
@@ I Introduction
Mathematical models grounded in physical laws offer profound insights into the behavior of complex systems across science and engineering. These models clarify the underlying mechanisms governing system dynamics and empower predictions and innovations in technology and natural science. Traditionally, model derivation has leaned heavily on theoretical and empirical knowledge, often requiring expert knowledge and intuition. Data-driven methods have become capable of assisting in developing mathematical models and constructing models that provide advanced predictions Brunton and Kutz (2022). These data-driven methods involve the sparse identification Brunton, Proctor, and Kutz (2016); de Silva et al. (2020); Kaheman, Brunton, and Nathan Kutz (2022); Fasel et al. (2022), symbolic regression Schmidt and Lipson (2009); Bongard and Lipson (2007); Reinbold et al. (2021); Udrescu and Tegmark (2020); Cranmer et al. (2020); Lemos et al. (2023), and physics-informed machine learning methods Karniadakis et al. (2021); Raissi, Perdikaris, and Karniadakis (2019); Jia et al. (2021); Rosofsky, Al Majed, and Huerta (2023). These methods have emerged as powerful tools for deriving governing equations directly from data, overcoming the limitations of conventional expert-dependent approaches.
Rheology is one of the scientific fields that address the properties of flowing matter, which plays a crucial role in many industries, such as designs of chemical processes, by providing insights into the flow behavior of complex fluids. One of the roles of rheology is to discover or derive governing equations that relate deformation and stress, referred to as constitutive equations Larson (1988). Accurate constitutive equations are necessary to predict the flows of complex fluids under complex boundary conditions. Derivations of constitutive equations starting from the principles of continuum mechanics have achieved significant success in the field of rheology, yielding several practical constitutive equations Larson (1988). Nevertheless, it is generally difficult to theoretically obtain constitutive equations for complex fluids from molecular models. Such cases usually explore mesoscopic coarse-grained models, which are based on molecular theories and suitable for numerical simulations. For example, for polymeric liquids, standard molecular theories have been proposed Rouse (1953); Doi and Edwards (1986) and refined mesoscopic models have been constructed based on them Masubuchi et al. (2001); Doi and Takimoto (2003); Likhtman (2005). In these models, the motion of individual (coarse-grained) molecules is numerically tracked. Although these models can reproduce rheological data with high accuracy, they require significantly more computational time compared to constitutive equations. Thus, a clear methodology is desired for obtaining constitutive equations from available data.
@@
Despite the success of NNs and GPRs, their black-box nature often obscures the underlying physics, making symbolic regression techniques more appealing for transparency and interoperability Brunton and Kutz (2022). For example, this technique can successfully identify the governing equations for fluid flows Reinbold et al. (2021). Moreover, the sparse identification of nonlinear dynamics (SINDy) Brunton, Proctor, and Kutz (2016), which is one of such methods, has been utilized to track (reduced order) dynamics in the field of fluid mechanics Fukami et al. (2021). Inspired by these successes, symbolic regression methods have recently started to be used in the field of rheology as well. For example, Mohammadamin and coworkers relied on SINDy for flexibly identifying the constitutive equations for an elasto-visco-plastic fluid Mahmoudabadbozchelou et al. (2024). Generally, the performance of SINDy is significantly influenced by the training data collection method, the candidate terms selected, and the optimization method. However, a comprehensive study to test SINDy for rheological data has not yet been conducted. Before applying SINDy to real-world rheological data, it is highly desirable to investigate fundamental learning strategies such as how to collect training data among various rheological tests and which optimization method to implement.
@@ II Methods
where ${\bm{t}}_{\mu\nu}$ ($\mu\nu\in\{xx,yy,zz,xy,yz,zx\}$) is the stress data for $n$ sequential times. In this study, the stress data are collected by applying $\bm{\kappa}(t)$ to systems of prescribed constitutive equations or mesoscopic models of viscoelastic fluids. The time derivatives of the stress data $\dot{\bm{T}}$ are computed by a numerical differentiation method. The remaining required data is the velocity gradient data $\bm{K}$ summarized as
@@
${\bm{\Theta}}=\left[\begin{array}[]{ccccccc}{\bm{1}}&{\bm{T}}&{\bm{K}}&({\bm{T}}\otimes{\bm{T}})&({\bm{T}}\otimes{\bm{K}})&({\bm{K}}\otimes{\bm{K}})&\cdots\end{array}\right],$ (6)
where ${\bm{T}}\otimes{\bm{K}}$, for example, denotes all possible combinations of the products of the row components in ${\bm{T}}$ and ${\bm{K}}$ for each time $t_{i}$ ($1\leq i\leq n$). The library ${\bm{\Theta}}$ can incorporate not only polynomials but also other functions, such as sinusoidal functions. If the library contains unnecessary functions or does not contain necessary functions, the correct expression cannot be obtained by Rheo-SINDy. Thus, functions to be included in the library must be selected carefully by utilizing expertise in rheology and knowledge of the target data. The detailed procedure varies from case to case and is provided in Sec. IV.
@@
$\bm{\Xi}=\left[\begin{array}[]{cccc}{\bm{\xi}}_{xx}&{\bm{\xi}}_{yy}&\cdots&{\bm{\xi}}_{zx}\\ \end{array}\right]=\left[\begin{array}[]{cccc}\xi_{xx,1}&\xi_{yy,1}&\cdots&\xi_{zx,1}\\ \xi_{xx,2}&\xi_{yy,2}&\cdots&\xi_{zx,2}\\ \vdots&\vdots&\ddots&\vdots\\ \xi_{xx,N_{\bm{\Theta}}}&\xi_{yy,N_{\bm{\Theta}}}&\cdots&\xi_{zx,N_{\bm{\Theta}}}\end{array}\right],$ (8)
where $N_{\bm{\Theta}}$ is the total number of library functions. To determine the coefficients $\bm{\Xi}$, we solve the following optimization problem for each row:
@@
and $R({\bm{\xi}}_{\mu\nu})$ is the regularization term. To obtain a sparse solution of $\bm{\Xi}$ from rheological data, we apply the following five methods Fukami et al. (2021): (i) the sequentially thresholded least square algorithm (STLSQ), (ii) sequentially thresholded Ridge regression (STRidge), (iii) least absolute shrinkage and selection operator (Lasso), (iv) Elastic-Net (E-Net), and (v) adaptive-Lasso (a-Lasso). These methods employ different regularization terms to obtain sparse solutions (see Sec. S1 in the supplementary material for details). Each method has a hyperparameter $\alpha$ to penalize the solution complexity, which is to be tuned for obtaining good predictive yet parsimonious representations. For this purpose, we test various $\alpha$ values and pick the one with the smallest number of terms among the results whose error has the same order as the minimum error (when $\alpha$ is sufficiently small).
@@ III Case Studies
For case studies, we first test whether Rheo-SINDy can find appropriate constitutive equations from training data generated by phenomenological constitutive equations, namely the Upper Convected Maxwell (UCM) model and the Giesekus model (the details are summarized in Sec. S2 in the supplementary material). Subsequently, we apply Rheo-SINDy to data generated by several dumbbell-based models. This section provides a brief overview of the dumbbell models used in this study and the conditions for creating the training datasets.

### v3

@@ Abstract
Rheology plays a pivotal role in understanding the flow behavior of fluids by discovering governing equations that relate deformation and stress, known as constitutive equations. Despite the importance of these equations, current methods for deriving them lack a systematic methodology, often relying on sense of physics and incurring substantial costs. To overcome this problem, we propose a novel method named Rheo-SINDy, which employs the sparse identification of nonlinear dynamics (SINDy) algorithm for discovering constitutive models from rheological data. Rheo-SINDy was applied to five distinct scenarios, four with well-established constitutive equations and one without predefined equations. Our results demonstrate that Rheo-SINDy successfully identified accurate models for the known constitutive equations and derived physically plausible approximate models for the scenario without established equations. Notably, the identified approximate models can accurately reproduce nonlinear shear rheological properties, especially at steady state, including shear thinning. These findings validate the robustness of Rheo-SINDy in handling data complexities and underscore its efficacy as a tool for advancing the development of data-driven approaches in rheology.
@@ I Introduction
Mathematical models grounded in physical laws offer profound insights into the behavior of complex systems across science and engineering. These models clarify the underlying mechanisms governing system dynamics and empower predictions and innovations in technology and natural science. Traditionally, model derivation has leaned heavily on theoretical and empirical knowledge, often requiring expert intuition. Data-driven methods have become capable of assisting in developing mathematical models and constructing models that provide advanced predictions Brunton and Kutz (2022). These data-driven methods involve the sparse identification Brunton, Proctor, and Kutz (2016); de Silva et al. (2020); Kaheman, Brunton, and Nathan Kutz (2022); Fasel et al. (2022), symbolic regression Schmidt and Lipson (2009); Bongard and Lipson (2007); Reinbold et al. (2021); Udrescu and Tegmark (2020); Cranmer et al. (2020); Lemos et al. (2023), and physics-informed machine learning methods Karniadakis et al. (2021); Raissi, Perdikaris, and Karniadakis (2019); Jia et al. (2021); Rosofsky, Al Majed, and Huerta (2023). These methods have emerged as powerful tools for deriving governing equations directly from data.
Rheology is one of the scientific fields that address the properties of flowing matter, which plays a crucial role in many industries, such as designs of chemical processes, by providing insights into the flow behavior of complex fluids. One of the roles of rheology is to discover or derive governing equations that relate deformation and stress, referred to as constitutive equations Larson (1988). Accurate constitutive equations are necessary to predict the flows of complex fluids under complex boundary conditions. Derivations of constitutive equations starting from the principles of continuum mechanics have achieved significant success in the field of rheology, yielding several practical constitutive equations Larson (1988). Additionally, Ilg and Kröger proposed the derivation of constitutive equations following the guidelines of nonequilibrium thermodynamics Ilg and Kröger (2011). Nevertheless, it is generally difficult to theoretically obtain constitutive equations for complex fluids from molecular models. Such cases usually explore mesoscopic coarse-grained models, which are based on molecular theories and suitable for numerical simulations. For example, for polymeric liquids, standard molecular theories have been proposed Bird, Armstrong, and Hassager (1987); Rouse (1953); Doi and Edwards (1986) and refined mesoscopic models have been constructed based on them Masubuchi et al. (2001); Doi and Takimoto (2003); Likhtman (2005). In these models, the motion of individual (coarse-grained) molecules is numerically tracked. Although these models can reproduce rheological data with high accuracy, they require significantly more computational time compared to constitutive equations. Thus, a clear methodology is desired for obtaining constitutive equations from available data with the assistance of the rheological knowledge.
@@
Despite the success of NNs and GPRs, with a few exceptions such as the recent work of Lennon and coworkers Lennon, McKinley, and Swan (2023), their black-box nature often obscures the underlying physics, making symbolic regression techniques more appealing for transparency and interoperability Brunton and Kutz (2022). For example, this technique can successfully identify the governing equations for fluid flows Reinbold et al. (2021). Moreover, the sparse identification of nonlinear dynamics (SINDy) Brunton, Proctor, and Kutz (2016), which is one of such methods, has been utilized to track (reduced order) dynamics in the field of fluid mechanics Fukami et al. (2021). Inspired by these successes, symbolic regression methods have recently started to be used in the field of rheology as well. For example, Mohammadamin and coworkers relied on SINDy for flexibly identifying the constitutive equations for an elasto-visco-plastic fluid Mahmoudabadbozchelou et al. (2024). Generally, the performance of SINDy is significantly influenced by the training data collection method, the candidate terms selected, and the optimization method. However, a comprehensive study to test SINDy for rheological data has not yet been conducted. Before applying SINDy to real-world rheological data, it is highly desirable to investigate fundamental learning strategies such as how to collect training data among various rheological tests and which optimization method to implement.
@@ II Methods
where ${\bm{t}}_{\mu\nu}$ ($\mu\nu\in\{xx,yy,zz,xy,yz,zx\}$) is the stress data for $n$ sequential times. In this study, the stress data are collected by applying $\bm{\kappa}(t)$ to systems of prescribed constitutive equations or mesoscopic models of viscoelastic fluids. The time derivatives of the stress data $\dot{\bm{T}}$ are computed by a numerical differentiation method. When applying Rheo-SINDy with experimental data, we need to address the errors in time derivative data due to limited experimental time resolutions. To address this, it is effective to use a high-accuracy numerical differentiation scheme, such as the differentiation scheme with total variation regularization Chartrand (2011), as reported by Brunton and coworkers Brunton, Proctor, and Kutz (2016). Nevertheless, in this study, we use a simple finite difference method and demonstrate that the correct equations can be regressed with this method. The remaining required data is the velocity gradient data $\bm{K}$ summarized as
@@
${\bm{\Theta}}$ $=\left[\begin{array}[]{cccc}{\bm{\theta}}_{1}&{\bm{\theta}}_{2}&\cdots&{\bm{\theta}}_{N_{\bm{\Theta}}}\\ \end{array}\right]$ (7)
$=\left[\begin{array}[]{ccccccc}{\bm{1}}&{\bm{T}}&{\bm{K}}&({\bm{T}}\otimes{\bm{T}})&({\bm{T}}\otimes{\bm{K}})&({\bm{K}}\otimes{\bm{K}})&\cdots\end{array}\right],$ (9)
where $N_{\bm{\Theta}}$ is the total number of library functions and ${\bm{T}}\otimes{\bm{K}}$, for example, denotes all possible combinations of the products of the row components in ${\bm{T}}$ and ${\bm{K}}$ for each time $t_{i}$ ($1\leq i\leq n$). The library ${\bm{\Theta}}$ can incorporate not only polynomials but also other functions, such as sinusoidal functions. If the library does not contain necessary functions, the correct expression cannot be obtained by Rheo-SINDy. Thus, functions to be included in the library must be selected carefully by utilizing expertise in rheology and knowledge of the target data. The detailed procedure varies from case to case and is provided in Sec. IV.
@@
$\bm{\Xi}=\left[\begin{array}[]{cccc}{\bm{\xi}}_{xx}&{\bm{\xi}}_{yy}&\cdots&{\bm{\xi}}_{zx}\\ \end{array}\right]=\left[\begin{array}[]{cccc}\xi_{xx,1}&\xi_{yy,1}&\cdots&\xi_{zx,1}\\ \xi_{xx,2}&\xi_{yy,2}&\cdots&\xi_{zx,2}\\ \vdots&\vdots&\ddots&\vdots\\ \xi_{xx,N_{\bm{\Theta}}}&\xi_{yy,N_{\bm{\Theta}}}&\cdots&\xi_{zx,N_{\bm{\Theta}}}\end{array}\right].$ (11)
To determine the coefficients $\bm{\Xi}$, we solve the following optimization problem for each row:
@@
and $R({\bm{\xi}}_{\mu\nu})$ is the regularization term. In SINDy, the appropriate optimization method generally depends on the specific problem, as has already been demonstrated by Fukami and coworkers Fukami et al. (2021). To obtain a sparse solution of $\bm{\Xi}$ from rheological data, we apply the following five methods Fukami et al. (2021): (i) the sequentially thresholded least square algorithm (STLSQ), (ii) sequentially thresholded Ridge regression (STRidge), (iii) least absolute shrinkage and selection operator (Lasso), (iv) Elastic-Net (E-Net), and (v) adaptive-Lasso (a-Lasso).
These methods employ different regularization terms as shown in Table 1.
Table 1: The regularization term $R({\bm{\xi}}_{\mu\nu})$ for the sparse regression methods.
method regularization term $R({\bm{\xi}}_{\mu\nu})$
STLSQ $\lambda_{0}||\bm{\xi}_{\mu\nu}||_{0}$
STRidge $\lambda_{0}||\bm{\xi}_{\mu\nu}||_{0}+\lambda_{2}||\bm{\xi}_{\mu\nu}||_{2}^{2}$
Lasso $\lambda_{1}||\bm{\xi}_{\mu\nu}||_{1}$
E-Net $\lambda_{1}||\bm{\xi}_{\mu\nu}||_{1}+\lambda_{2}||\bm{\xi}_{\mu\nu}||_{2}^{2}$
a-Lasso $\lambda_{1}||\bm{\xi}_{\mu\nu}^{\prime}||_{1}$
The hyperparameters of $\ell_{i}$ norm ($i=0,1,2$) are denoted as $\lambda_{i}$ ($>0$). The $\ell_{0}$ and $\ell_{1}$ norms are defined as
$||\bm{\xi}_{\mu\nu}||_{0}=\sum_{j}\delta(\xi_{\mu\nu,j})$ (14)
and
$||\bm{\xi}_{\mu\nu}||_{1}=\sum_{j}|\xi_{\mu\nu,j}|,$ (15)
where $\delta(\xi_{\mu\nu,j})$ is the Kronecker delta function, which is equal to $1$ if $\xi_{\mu\nu,j}\neq 0$ and 00 otherwise. The vector $\bm{\xi}^{\prime}_{\mu\nu}$ in the a-Lasso is defined as $\bm{\xi}^{\prime}_{\mu\nu}=\bm{w}_{\mu\nu}\otimes\bm{\xi}_{\mu\nu}$, where $\otimes$ is the element-wise product and $\bm{w}_{\mu\nu}$ is the adaptive weight vector and its $j$-th element is defined as $w_{\mu\nu,j}=|\xi_{\mu\nu,j}|^{-\delta}$ with $\delta$ being the positive constant.
The STLSQ and STRidge were implemented by iteratively conducting the least square regression and the Ridge regression, respectively, while setting the coefficients with smaller absolute values than a certain threshold $\alpha$ ($>0$) to zero based on the original papers Brunton, Proctor, and Kutz (2016); Rudy et al. (2017). Since this coefficient selection by $\alpha$ replaces the role of the $\ell_{0}$ norm, $\lambda_{0}$ is set to zero in this implementation. In the STRidge, the hyperparameter $\lambda_{2}$ was set to $0.05$. The Lasso, E-Net, and a-Lasso were implemented using the scikit-learn library Pedregosa et al. (2011). In this library, the loss functions $L({\bm{\xi}}_{\mu\nu})$ for the Lasso and E-Net are respectively defined as
$L({\bm{\xi}}_{\mu\nu})=\frac{1}{2n}||\dot{\bm{t}}_{\mu\nu}-{\bm{\Theta}}\bm{\xi}_{\mu\nu}||_{2}^{2}+\alpha||\bm{\xi}_{\mu\nu}||_{1}$ (16)
and
$L({\bm{\xi}}_{\mu\nu})=\frac{1}{2n}||\dot{\bm{t}}_{\mu\nu}-{\bm{\Theta}}\bm{\xi}_{\mu\nu}||_{2}^{2}+\alpha\beta||\bm{\xi}_{\mu\nu}||_{1}+\frac{\alpha(1-\beta)}{2}||\bm{\xi}_{\mu\nu}||_{2}^{2},$ (17)
where $\beta$ is the $\ell_{1}$ ratio and $\alpha$ and $\beta$ are the hyperparameters. These two loss functions have the same form when $\beta=1$. In this study, $\beta$ for the E-Net was set to $0.5$. According to the original paper of the a-Lasso Zou (2006), it can be implemented as the Lasso problem as the following steps:
  1. 1.
    Define ${\bm{\Theta}}^{\prime}=\left[{\bm{\theta}}_{1}^{\prime},\cdots,{\bm{\theta}}_{N_{\bm{\Theta}}}^{\prime}\right]$, where ${\theta}_{j}^{\prime}={\theta}_{j}/w_{\mu\nu,j}$ ($j=1,\ldots,N_{\bm{\Theta}}$).
  2. 2.
    Solve the Lasso problem to obtain $\hat{\bm{\xi}}^{\prime}_{\mu\nu}$ using Eq. (16) with ${\bm{\Theta}}^{\prime}$.
  3. 3.
    Output $\hat{\xi}_{\mu\nu,j}=\hat{\xi}^{\prime}_{\mu\nu,j}/w_{\mu\nu,j}$ ($j=1,\ldots,N_{\bm{\Theta}}$)
The adaptive weight $w_{\mu\nu,j}$ depends on the coefficients, and thereby the output coefficients can be varied in each iteration. To obtain the converged solution, we initialized the weights as unit vectors $\bm{w}=\bm{1}$ and repeated the above steps until the coefficients $\hat{\xi}_{\mu\nu,j}$ no longer change Fukami et al. (2021). Here, the hyperparameter $\delta$ was set to $3$ (see Sec. S1 in the supplementary material for the effect of $\delta$). As shown above, each method has a hyperparameter $\alpha$ to penalize the solution complexity, which is to be tuned for obtaining good predictive yet parsimonious representations. For this purpose, we test various $\alpha$ values and pick the one with the smallest number of terms among the results whose error has the same order as the minimum error (when $\alpha$ is sufficiently small).
@@ III Case Studies
For case studies, we first test whether Rheo-SINDy can find appropriate constitutive equations from training data generated by phenomenological constitutive equations, namely the Upper Convected Maxwell (UCM) model and the Giesekus model. Subsequently, we apply Rheo-SINDy to data generated by several dumbbell-based models. This section provides a brief overview of the UCM, Giesekus, and dumbbell models used in this study and the conditions for creating the training datasets. Table 2 summarizes the conditions for generating training datasets.
Table 2: Conditions for generating training datasets.
model parameter shear type111“S” means the steady shear flow and “O” means the oscillatory shear flow. shear parameters $\Delta t$ $\Delta t_{\mathrm{train}}$ simulation time
UCM S $\dot{\gamma}\in\{1,1.7,\ldots,100\}$ [^2] Ten equally spaced values on a logarithmic scale between $\dot{\gamma}=1$ and $\dot{\gamma}=100$ were used. $10^{-4}$ $10^{-2}$ $10$
UCM O $\gamma_{0}=2$, $\omega=1$ $10^{-4}$ $10^{-2}$ $100$
Giesekus $\alpha_{\rm G}=0.5$ O $\gamma_{0}=2$, $\omega\in\{0.1,0.2,\ldots,1\}$ $10^{-4}$ $10^{-2}$ $100$
Hookean dumbbell [^3] Three different numbers of dumbbells ($N_{\rm p}\in\{10^{3},10^{4},10^{5}\}$) were addressed. $n_{\mathrm{K}}=10$ O $\gamma_{0}=2$, $\omega=0.5$ $10^{-3}$ $10^{-2}$ $100$
FENE-P dumbbell [^4] The closed form expression ($N_{\rm p}\to\infty$) was examined. $n_{\mathrm{K}}=10$ O $\gamma_{0}=2$ or $8$, $\omega\in\{0.1,0.2,\ldots,1\}$ $10^{-4}$ $10^{-2}$ $100$
FENE dumbbell [^5] The number of dumbbells was set to $N_{\rm p}=10^{4}$. $n_{\mathrm{K}}=10$ O $\gamma_{0}=2$ or $8$, $\omega\in\{0.1,0.2,\ldots,1\}$ $10^{-4}$ $10^{-2}$ $100$
III.1 Upper Convected Maxwell (UCM) Model
The simplest constitutive equation for viscoelastic fluids is the UCM model Larson (1988) shown as
$\frac{{\rm d}{\bm{\tau}}}{{\rm d}t}-{\bm{\tau}}\cdot{\bm{\kappa}}^{\rm T}-{\bm{\kappa}}\cdot{\bm{\tau}}=-\frac{1}{\lambda}\bm{\tau}+2G\bm{D}.$ (18)
Here, the left-hand side of Eq. (18) is the upper-convected time derivative of $\bm{\tau}$, $\bm{\kappa}^{\rm T}$ is the transposed $\bm{\kappa}$, $\lambda$ is the relaxation time, $G$ is the modulus, and $\bm{D}$ is the deformation rate tensor defined as $\bm{D}=(\bm{\kappa}+{\bm{\kappa}}^{\rm T})/2$. Using $\lambda$ as the unit time and $G$ as the unit stress (i.e., $\lambda=G=1$), we can obtain dimensionless expressions for time $\tilde{t}=t/\lambda$, velocity gradient tensor $\tilde{\bm{\kappa}}=\lambda\bm{\kappa}$, and stress $\tilde{\bm{\tau}}={\bm{\tau}}/G$. In what follows, we omit the tilde in dimensionless variables for simplicity. The dimensionless form of the UCM model under shear flow is thus written as
${\dot{\tau}}_{xx}=-\tau_{xx}+2\tau_{xy}\kappa_{xy},$ (19)
${\dot{\tau}}_{yy/zz}=-{\tau}_{yy/zz}=0,$ (20)
Here, since the initial conditions for $\bm{\tau}$ are set to the values of $\bm{\tau}$ at equilibrium, namely $\bm{\tau}=\bm{0}$, ${\tau}_{yy/zz}$ of the UCM model is zero under shear flow.
For the UCM model, we generate training data by numerically solving Eqs. (19)–(21) under two shear flow scenarios: steady shear and oscillatory shear tests. For the steady shear test, the shear rate is kept constant ($\kappa_{xy}=\dot{\gamma}$) across various values ($\dot{\gamma}\in\{1,1.7,2.8,4.6,7.7,13,22,36,60,100\}$) with simulations running from $t=0$ to $t=10$ using a time step of $\Delta t=1.0\times 10^{-4}$. The oscillatory shear test introduces a time-dependent oscillatory shear strain, $\gamma(t)=\gamma_{0}\sin(\omega t)$ (i.e., $\kappa_{xy}(t)=\dot{\gamma}(t)=\gamma_{0}\omega\cos(\omega t)$), with $\gamma_{0}=2$ and $\omega=1$, over a period from $t=0$ to $t=100$, employing the same time step. In both tests, data are collected at intervals of $\Delta t_{\rm train}=1\times 10^{-2}$, resulting in a total of $10^{4}$ data points for the training data.
III.2 Giesekus Model
The Giesekus model, which is one of the most popular phenomenological constitutive equations Giesekus (1982), shows typical shear rheological properties and is used to fit various complex fluids, including polymer solutions and wormlike micellar solutions. The tensorial form of the Giesekus constitutive equation can be written as
$\frac{{\rm d}{\bm{\tau}}}{{\rm d}t}-{\bm{\tau}}\cdot{\bm{\kappa}}^{\rm T}-{\bm{\kappa}}\cdot{\bm{\tau}}=-\frac{1}{\lambda}\bm{\tau}-\frac{\alpha_{\rm G}}{G\lambda}\bm{\tau}\cdot\bm{\tau}+2G\bm{D},$ (22)
where $\alpha_{\rm G}$ is the parameter governing the nonlinear response of the Giesekus model. The Giesekus equation under shear flow is thus given by
${\dot{\tau}}_{xx}$ $=$ $-\tau_{xx}-\alpha_{\rm G}(\tau_{xx}^{2}+\tau_{xy}^{2})+2\tau_{xy}\kappa_{xy},$ (23)
${\dot{\tau}}_{yy}$ $=$ $-\tau_{yy}-\alpha_{\rm G}(\tau_{yy}^{2}+\tau_{xy}^{2}),$ (24)
${\dot{\tau}}_{zz}$ $=$ $0,$ (25)
${\dot{\tau}}_{xy}$ $=$ $-\tau_{xy}-\alpha_{\rm G}(\tau_{xx}+\tau_{yy})\tau_{xy}+\tau_{yy}\kappa_{xy}+\kappa_{xy}.$ (26)
Here, all quantities are non-dimensionalized by using $\lambda$ as the unit time and $G$ as the unit stress. From Eqs. (23)–(26), the total number of collect terms in the Giesekus model is $12$.
We generate the training data by solving Eqs. (23)–(26) numerically with $\alpha_{\rm G}=0.5$ and $\Delta t=1\times 10^{-4}$. We note that the Giesekus model with $\alpha_{\rm G}=0.5$ gives sufficient nonlinear features under shear flow. We applied the oscillatory shear flow with $\gamma_{0}=2$ and various $\omega$ values ($\omega\in\{0.1,0.2,\ldots,1\}$) for $0\leq t\leq 100$. From the computed stress data, we collected data at the interval of $\Delta t_{\rm train}=1\times 10^{-2}$.
III.3 Dumbbell Models